• Find the volume of the solid generated by revolving about the x-axis bounded by the curve, y2 = 9x and the line y = 3x
1
Expert's answer
2021-06-14T15:36:08-0400

Let us find the volume of the solid generated by revolving about the x-axis bounded by the curve, "y^2 = 9x" and the line "y = 3x." Firstly, let us find the points of intersection of the curve and the line: "(3x)^2=9x" implies "9x^2=9x," and hence "x_1=0" and "x_2=1." Let us use the formula "V=\pi\int_{x_1}^{x_2}(y_1^2(x)-y^2_2(x))dx."


"V=\pi\int_0^1(9x-(3x)^2)dx=9\pi\int_0^1(x-x^2)dx=9\pi(\frac{x^2}{2}-\frac{x^3}{3})|_0^1=
9\pi(\frac{1}{2}-\frac{1}{3})=9\pi\cdot\frac{1}{6}=\frac{3}{2}\pi."